# AMBER - Advanced SegFormer for Multi-Band Image Segmentation: an application to Hyperspectral Imaging

### v1

@@ Abstract
Deep learning has revolutionized the field of hyperspectral image (HSI) analysis, enabling the extraction of complex and hierarchical features. While convolutional neural networks (CNNs) have been the backbone of HSI classification, their limitations in capturing global contextual features have led to the exploration of Vision Transformers (ViTs). This paper introduces AMBER, an advanced SegFormer specifically designed for multi-band image segmentation. AMBER enhances the original SegFormer by incorporating three-dimensional convolutions to handle hyperspectral data. Our experiments, conducted on the Indian Pines, Pavia University, and PRISMA [^1] Project carried out using PRISMA Products, © of the Italian Space Agency (ASI), delivered under an ASI License to use. datasets, show that AMBER outperforms traditional CNN-based methods in terms of Overall Accuracy, Kappa coefficient, and Average Accuracy on the first two datasets, and achieves state-of-the-art performance on the PRISMA dataset.
@@
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
@@ 1 Introduction
In recent years, deep learning algorithms, particularly convolutional neural networks (CNNs), have become prominent in HSI classification due to their ability to learn hierarchical features from raw data. CNN-based models have shown considerable success in extracting spectral-spatial features. However, the inherent limitation of CNNs in capturing spectral-spatial long-range dependencies has spurred interest in transformer-based models.
Vision Transformers (ViTs)[1], with their self-attention mechanism, have demonstrated remarkable success in various image-processing tasks. They are adept at capturing global contextual information, which is crucial for HSI classification. Many deep learning models have adopted the transformer architecture for hyperspectral data, and have achieved good results.
This paper presents AMBER, an advanced version of SegFormer [2] tailored for hyperspectral image segmentation. By integrating three-dimensional convolutions, AMBER effectively processes multi-band images, preserving spatial dimensions and enhancing feature extraction. We evaluate AMBER on the Indian Pines, Pavia University, and PRISMA datasets, demonstrating its superior performance compared to traditional methods.
@@ 3 Methodology
This section introduces AMBER, a new and advanced SegFormer for Multi-Band Image Segmentation. As the original SegFormer [2], AMBER consists of two main modules: a hierarchical Transformer encoder which generates high-resolution coarse features and low-resolution fine features; and (2) a lightweight All-MLP decoder used to fuse multi-level features together. The final layer of the decoder performs dimensionality reduction to generate a bi-dimensional semantic segmentation mask from a three-dimensional input image. Unlike the original SegFormer, AMBER is suited to cope with multi-band images thanks to its three dimensional convolution. Moreover, AMBER preserves the input spatial dimensions $H$ and $W$: this in order to maximize accuracy at the price of alight increase in the number of trainable parameters. For an overview of the architecture, see Fig. 1.
Figure 1: The proposed AMBER framework consists of two main modules: A hierarchical Transformer encoder to extract spatial/spectral features; and a All-MLP decoder to fuse these multi-level features, reduce the spectral dimension, and predict the semantic segmentation mask.
Given an image of $D\times H\times W$ (where $D$ is the spectral dimension), we first divide it into patches of $3\times 3\times 3$ using a three-dimensional convolution layer (the patch size is the kernel dimension of the convolutional layer). We then use these patches as input to the hierarchical Transformer encoder to obtain multi-level features at 1/4, 1/8, 1/16 of the original image resolution. We then pass these multi-level features to the All-MLP decoder to predict the segmentation mask at a $\frac{D}{4}\times\frac{H}{4}\times\frac{W}{4}\times N_{cls}$ resolution, where $N_{cls}$ is the number of classes. To provide a detailed explanation of our model architecture, we will follow the same steps as in the original paper [2].
@@ 3.1 Hierarchical Transformer Encoder
Given an input image with a resolution of $D\times H\times W$, we perform patch merging to obtain a hierarchical feature map $F_{i}$ with a resolution of $\frac{Dqw}{2^{i+1}}\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C_{i}$, where $i\in{1,2,3,4}$, and $C_{i+1}$ is larger than $C_{i}$.
@@
$\text{Attention}(Q,K,V)=\text{Softmax}(\frac{QK^{T}}{\sqrt{d_{head}}})V.$ (1)
@@ 3.2 Lightweight All-MLP Decoder
The All-MLP decoder consists of five main steps. First, multi-level features from the MiT encoder go through a MLP layer to unify the channel dimension. Then, in a second step, features are up-sampled to the original three-dimensional image and concatenated together. Third, a MLP layer is adopted to fuse the concatenated features. Fourth, a convolutional layer is used with a kernel $(D\times 1\times 1)$ to reduce the image dimension to a bi-dimensional prediction. Finally, a two-dimensional convolutional layer with a kernel size of 1 is used to predict the segmentation mask, which has a $H\times W\times N_{cls}$ resolution, where $N_{cls}$ is the number of classes.
4 Experimental and analysis
@@ 4.1 Data and experimental environment description
In order to test the proposed architecture, we utilize hyperspectral imaging, which records a detailed spectrum of light for each pixel and provides an invaluable source of information regarding the physical nature of the different materials, leading to the potential of a more accurate classification. Experiments were conducted firstly, on two of the most widely used public dataset - Indian Pines [13] and Pavia University [13] hyperspectral datasets - and, finally, on a hyperspectral image acquired from the PRISMA satellite. Here is an overview of the datasets.
Indian Pines and Pavia University datasets
The Indian Pines and Pavia University data sets consists of images with the absorption band caused by water mist removed. The Indian Pines image is obtained from an AVIRIS spectrometer in a forest area in Northwestern Indiana, USA. The image size is $145\times 145$ pixels, the wavelength is $0.40-2.50\,\mu$m, and the number of bands is $200$, the spatial resolution is $20$ m. The Pavia University image is obtained by ROSIS spectrometer at the University of Pavia, Italy. The image size is $610\times 340$ pixels, the wavelength is $0.43-0.86\,\mu$m, the number of bands is $103$, and the spatial resolution is $1.3$ m. The two datasets are divided into 16 categories and 9 categories respectively. Their false-color images and ground-truth images are shown in Figs. 2 and 3. The ground-truth classes and their respective samples number are shown in Table 1 and Table 2
@@ 4.2 Implementation details
We trained AMBER and made inferences using the IBiSCo Data Center (Infrastructure for Big Data and Scientific Computing). Information about the Data Center is available at https://ibiscohpc-wiki.scope.unina.it/. Specifically, AMBER was trained using four Tesla V100. Unlike the original SegFormer[2], training was performed from scratch without pre-training the encoder and randomly initializing the decoder. During the training phase, we applied random flipping to the patches and excluded the undefined class from the computation of the loss function. We set crop size to $D\times 32\times 32$ (where $D$ is the spectral dimension of the datasets) on all the three datasets. We trained the model using an SGD optimizer for all experiments. We used a batch size of 3 for experiments on the Indian Pines dataset, 10 for those on the Pavia University dataset, and 4 for those on the PRISMA dataset. In all of our experiments, we used the Categorical Cross Entropy loss function. The learning rate was set to 0.01 for all the experiments, and it remained constant throughout the training process. We used the following numbers of epochs for experiments on three datasets: 50 for Indian Pines, 30 for Pavia University, and 50 for PRISMA. We report semantic segmentation performance using Overall Accuracy (OA), Kappa Coefficient (Kappa), and Average Accuracy (AA). These metrics will be explained in the following section.
@@ 4.4 Data Preprocessing
We constructed the datasets using only random patches, unlike other papers that employ various data balancing strategies (for example, see [15]). The purpose was both to assess AMBER’s ability to learn from an unbalanced dataset, and to create a dataset with a consistent number of patches. The center of each patch is a pixel randomly selected from pixels that are not located on the edges of the images. For the Pavia University dataset and the Indian Pines dataset, the center is a pixel that does not belong to the undefined class to avoid building patches with only undefined pixels. After creating the dataset, we addressed the well-known issue of pixel overlap between the training and test sets (effect of the random patches strategy), which can lead to information leakage [16, 17]. To mitigate this, we designated 20% of the patches for training and 80% for testing for the Indian Pines and Pavia University datasets, and 10% for training and 90% for testing for the PRISMA dataset.
Before the training phase, it is common to perform dimensionality reduction on the spectral dimension of hyperspectral images (e.g., PCA, SVD, or Gabor Filter [18, 19]) to address information redundancy. Conversely, our strategy was to keep the patches with the full spectral dimension, favoring a fine-grained classification, and leaving the task of extracting relevant features to AMBER’s encoder.
@@ 5.1 Classification results
To evaluate the effectiveness of this method in classifying hyperspectral remote sensing images, we made a comparison with four mainstream methods: SVD/Unet [19], HSI-CNN [20], 3D-CNN [21] and contextual deep CNN [22]. It’s important to note that not all experiments involve pixel-by-pixel classification (or semantic segmentation). Only SVD/Unet, contextual deep CNN, and AMBER use this approach. On the other hand, HSI-CNN and 3D-CNN perform pure image classification, where the classes are determined by the centroid of the patch. This results in an increase in the metrics. Despite this, AMBER shows better results compared to the other methods on the Indian Pines dataset and Pavia University dataset and comparable results on the PRISMA dataset.
@@
The 3D-CNN architecture is a Multiscale 3D deep convolutional neural network (M3D-DCNN) which could jointly learn both 2D Multi-scale spatial feature and 1D spectral feature from HSI data in an end-to-end approach.
@@
Every neural network mentioned above was trained on the same patches using the configuration described in the corresponding papers [19] [20] [21] [22]. “Same patches” in this context refers to patches that have identical centroids. AMBER was trained on patches of dimension $D\times 32\times 32$ where D is the spectral dimension, as previously mentioned. SVD/UNet was trained on patches of $3\times 32\times 32$ where 3 is the number of singular values. HSI-CNN was trained using patches of $D\times 3\times 3$, 3D-CNN using $D\times 7\times 7$ patches, and contextual deep CNN using $D\times 5\times 5$ patches (D is, again, the spectral dimension). For every method, we conducted 5 experiments for each dataset. The classification accuracy of different methods on the three data sets is presented in Tables 4, 5, and 6, with the highest values highlighted in bold font. A comprehensive comparative analysis is presented to draw conclusions.
@@
The Overall Accuracy (OA), Kappa coefficients, and Average Accuracy (AA) of the AMBER method, calculated across 5 different experiments, are higher than those of the compared methods on both the Indian Pines dataset and the Pavia University dataset. On the Indian Pines dataset, the method achieved an OA of 99.74%, a Kappa coefficient of 0.9968, and an AA of 99.18%, representing a significant improvement compared to the other methods. In particular, AMBER demonstrates the ability to detect classes with few representations, such as class 7 (Grass-pasture-mowed), which has only 28 pixels. Other methods, such as SVD/Unet or HSI-CNN, struggle to detect this class, resulting in an accuracy per class of 0% . Another important feature is that, compared to other methods (especially SVD/Unet), AMBER has a very low standard deviation for OA, Kappa coefficient, and AA. This indicates a good ability to generalize: the results do not seem to depend on how the training data are chosen.
For the Pavia University dataset, the OA reached 99.94%, the Kappa coefficient reached 0.9991 and the AA reached 99.82%, mantaining an high accuracy on all the classes. As the Indian Pines dataset, AMBER has a very low standard deviation for OA, Kappa coefficient, and AA.
For the PRISMA dataset, the AMBER model achieved an Overall Accuracy of 90.90%, a Kappa coefficient of 0.8741, and an Average Accuracy of 90.92. While this result is slightly lower than the SVD/Unet method, it outperformed the other methods. Furthermore, the lower performances of AMBER on the PRISMA dataset are under investigation. AMBER is also able to improve the classification accuracy of a variety of features, such as Alfalfa, Grass-pasture-mowed and Soybean-mintill in the Indian Pines dataset; Asphalt, Gravel and Bitumen in the Pavia University dataset; mining area and build-up in the PRISMA dataset. This method is also able to perform segmentation on different types of hyperspectral images, maintaining high accuracy on both airborne (Pavia University and Indian Pines) and spaceborne (PRISMA) hyperspectral images.
@@
The Fig. 5, 6, and 7 show the false color image (a), the ground truth sample plots (b), and the AMBER classification result plots (c) and (d). Specifically, plot (c) is the predicted result (d) with an additional mask of the ground truth undefined pixels applied a posteriori to visually justify the high accuracy for all types of features presented in Tables 4, 5 and 6. Specifically, on the Pavia University dataset and the Indian Pines dataset, the ground truth (b) and the prediction (c) are nearly identical, as expected.
@@
Fig. 8 shows that AMBER accurately predicts the pixels classified as undefined in the ground-truth map (which are not used in the loss function computation). The figure shows an example of how AMBER can also segment complex structures. This particular structure belongs to the asphalt class in the Pavia University dataset.

### v2

@@ Abstract
Deep learning has revolutionized the field of hyperspectral image (HSI) analysis, enabling the extraction of complex spectral and spatial features. While convolutional neural networks (CNNs) have been the backbone of HSI classification, their limitations in capturing global contextual features have led to the exploration of Vision Transformers (ViTs). This paper introduces AMBER, an advanced SegFormer specifically designed for multi-band image segmentation. AMBER enhances the original SegFormer by incorporating three-dimensional convolutions, custom kernel sizes and a Funnelizer layer. This architecture enables to process hyperspectral data directly, without requiring spectral dimensionality reduction during preprocessing. Our experiments, conducted on three benchmark datasets (Salinas, Indian Pines and Pavia University) and on a dataset from the PRISMA [^1] Project carried out using PRISMA Products, © of the Italian Space Agency (ASI), delivered under an ASI License to use. satellite, show that AMBER outperforms traditional CNN-based methods in terms of Overall Accuracy, Kappa coefficient, and Average Accuracy on the first three datasets, and achieves state-of-the-art performance on the PRISMA dataset. These findings highlight AMBER’s robustness, adaptability to both airborne and spaceborne data, and its potential as a powerful solution for remote sensing and other domains requiring advanced analysis of high-dimensional data.
@@ 1 Introduction
In recent years, deep learning algorithms, particularly convolutional neural networks (CNNs), have become prominent in HSI classification due to their ability to learn hierarchical features from raw data. CNN-based models have shown considerable success in extracting spectral-spatial features [1, 2]. However, the inherent limitation of CNNs in capturing spectral-spatial long-range dependencies has spurred interest in transformer-based models [3].
Vision Transformers (ViTs) [4], with their self-attention mechanism, have demonstrated remarkable success in various image-processing tasks. They are adept at capturing global contextual information, which is crucial for HSI classification. Many deep learning models have adopted the transformer architecture for hyperspectral data, and have achieved good results (e.g., [5, 6, 7, 8, 9, 10, 11]).
This paper introduces AMBER, an advanced version of SegFormer [12], specifically designed for multi-band image segmentation. AMBER incorporates several key innovations: it integrates three-dimensional convolutions, employs custom kernel sizes and strides to preserve output spatial dimensions. It also includes a Funnelizer layer to reduce spectral dimensions for generating two-dimensional masks. These enhancements enable AMBER to effectively process both spectral and spatial dimensions without sacrificing critical information.
We evaluated AMBER on four datasets—Salinas, Indian Pines, Pavia University, and PRISMA—where it demonstrated superior performance compared to traditional methods on the Salinas, Indian Pines and Pavia University datasets, achieving state-of-the-art results on the PRISMA dataset. Notably, AMBER exhibits robustness across diverse scenarios and conditions, performing well with data from both airborne spectrometers (AVIRIS for Salinas and Indian Pines; ROSIS for Pavia University) and spaceborne spectrometers (PRISMA).
A distinctive feature of AMBER is its ability to handle hyperspectral data without requiring a preprocessing step to reduce spectral dimensions (e.g., SVD, or Gabor Filter [13, 14]), adopting a full-spectrum approach instead. This not only streamlines the workflow but also preserves essential spectral information, which is often crucial for accurate classification and analysis. The preservation of such spectral details is particularly beneficial for applications like land cover classification, mineral identification, and environmental monitoring, where subtle differences in spectral signatures can reveal significant insights.
AMBER’s innovative architecture significantly enhances representation learning by effectively addressing the unique challenges of hyperspectral data and, more broadly, complex multi-dimensional data cubes. Its ability to process such data makes it highly versatile, with potential applications extending beyond hyperspectral imaging. Future work will aim to broaden its applicability to additional domains, including medical imaging and astronomical data, where the analysis of multi-dimensional datasets plays a crucial role.
@@ 3 Methodology
This section introduces AMBER, a new and advanced SegFormer designed specifically for Multi-Band Image Segmentation. As the original SegFormer [12], AMBER consists of two main modules: a hierarchical Transformer encoder, which generates high-resolution coarse features and low-resolution fine features; and a lightweight All-MLP decoder, used to fuse multi-level features together.
Unlike the original SegFormer, AMBER is tailored to handle multi-band images through the integration of three-dimensional convolutions. Furthermore, AMBER preserves the input spatial dimensions $H$ and $W$ by employing custom kernel sizes and strides, ensuring higher accuracy. This enhancement comes with only a slight increase in the number of trainable parameters. A key addition in AMBER is the Funnelizer layer, which reduces the spectral dimensionality, enabling the generation of a two-dimensional segmentation mask from a three-dimensional input image. For an overview of the architecture, see Fig. 1.
Figure 1: The proposed AMBER framework consists of two main modules: A hierarchical Transformer encoder to extract spatial/spectral features; and a All-MLP decoder to fuse these multi-level features; the Funnelizer layer reduce the spectral dimension and predict the segmentation mask.
Given an image of $D\times H\times W$ (where $D$ represents the spectral dimension, $H$ denotes the height, and $W$ represents the width of the image), we first divide it into patches of $3\times 3\times 3$ using a three-dimensional convolution layer (the patch size is the kernel dimension of the convolutional layer). We then use these patches as input to the hierarchical Transformer encoder to obtain multi-level features at 1/4, 1/8, 1/16 of the original image resolution. We then pass these multi-level features to the All-MLP decoder to predict the segmentation mask at a $\frac{D}{4}\times\frac{H}{4}\times\frac{W}{4}\times N_{cls}$ resolution, where $N_{cls}$ is the number of classes. To provide a detailed explanation of our model architecture, we will follow almost the same steps as in the original paper [12].
@@ 3.1 Hierarchical Transformer Encoder
Given an input image with a resolution of $D\times H\times W$, we perform patch merging to obtain a hierarchical feature map $F_{i}$ with a resolution of ${\frac{D}{2^{i+1}}\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C_{i}}$, where $i\in{1,2,3,4}$, and $C_{i+1}$ is larger than $C_{i}$.
@@
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{head}}}\right)V.$ (1)
@@ 3.2 Lightweight All-MLP Decoder
The All-MLP decoder consists of five main steps. First, multi-level features from the MiT encoder go through a MLP layer to unify the channel dimension. Then, in a second step, features are up-sampled to the original three-dimensional image and concatenated together. Third, a MLP layer is adopted to fuse the concatenated features.
The Funnelizer layer, a convolution with a kernel size of $(D\times 1\times 1)$ reduces the spectral dimension, transforming the data into a two-dimensional representation. Finally, a two-dimensional convolutional layer with a kernel size of 1 is used to predict the segmentation mask, which has a resolution of $H\times W\times N_{cls}$, where $N_{cls}$ is the number of classes.
4 Experiments and Analysis
@@ 4.1 Data and experimental environment description
In order to evaluate the proposed architecture, we utilize hyperspectral imaging, which records a detailed spectrum of light for each pixel and provides an invaluable source of information regarding the physical nature of the different materials, leading to the potential of a more accurate classification. Experiments were conducted firstly, on three of the most widely used public datasets — Salinas, Indian Pines, and Pavia University [24]— as well as on a hyperspectral image acquired from the PRISMA satellite. Here is an overview of the datasets.
Salinas, Indian Pines and Pavia University datasets
The Salinas, Indian Pines, and Pavia University datasets consist of hyperspectral images with water absorption bands removed for better spectral analysis.
The Salinas dataset was acquired using the AVIRIS spectrometer over the Salinas Valley, California. It is characterized by high spatial resolution (3.7-meter pixels). After the removal of water absorption bands, the image contains 204 spectral bands and has dimensions of $512\times 217$ pixels.
The Indian Pines dataset was also captured by the AVIRIS spectrometer, covering a mixed agricultural and forested area in Northwestern Indiana, USA. The image has dimensions of $145\times 145$ pixels, a spectral range of $0.40-2.50,\mu$m, 200 spectral bands, and a spatial resolution of 20 meters.
The Pavia University dataset was acquired using the ROSIS spectrometer during a flight campaign over the University of Pavia, Italy. This dataset features an image with dimensions of $610\times 340$ pixels, a spectral range of $0.43-0.86,\mu$m, 103 spectral bands, and a spatial resolution of 1.3 meters.
The Salinas and Indian Pines datasets each contain 16 ground-truth classes, while the Pavia University dataset has 9 ground-truth classes. The false-color images and ground-truth maps for these datasets are shown in Figs. 2, 3, and 4, with the class labels and their respective sample counts provided in Tables 1, 2 and 3.
Figure 2: Salinas dataset. (a) False-color map; (b) Ground-truth map.
Table 1: Groundtruth classes for the Salinas scene and their respective samples number.
# Class Samples
1 Brocoli green weeds 1 2009
2 Brocoli green weeds 2 3726
3 Fallow 1976
4 Fallow rough plow 1394
5 Fallow smooth 2678
6 Stubble 3959
7 Celery 3579
8 Grapes untrained 11271
9 Soil vinyard develop 6203
10 Corn senesced green weeds 3278
11 Lettuce romaine 4wk 1068
12 Lettuce romaine 5wk 1927
13 Lettuce romaine 6wk 916
14 Lettuce romaine 7wk 1070
15 Vinyard untrained 7268
16 Vinyard vertical trellis 1807
[…]
@@ 4.2 Implementation details
We trained AMBER and conducted inferences using the IBiSCo Data Center (Infrastructure for Big Data and Scientific Computing). Detailed information about the Data Center can be found at https://ibiscohpc-wiki.scope.unina.it/. Specifically, AMBER was trained using four Tesla V100 GPUs, each providing 32 GB of memory. Training times varied based on dataset size and complexity: approximately 6 hours for the Salinas dataset, 5 hours for the Indian Pines dataset, and around 7 hours for both the Pavia University and PRISMA datasets. Unlike the original SegFormer [12], AMBER was trained entirely from scratch, with no pre-training for the encoder and random initialization of the decoder.
During the training phase, we applied random flipping to the patches and excluded the undefined class from the computation of the loss function. The crop size was set to $D\times 32\times 32$ (where $D$ is the spectral dimension of the datasets) for all four datasets. This crop size was chosen after empirical evaluation: larger patches led to poorer predictions as AMBER struggled to capture finer details, while smaller patches resulted in a loss of context, degrading performance as well.
We trained the model using an SGD optimizer for all experiments, with a batch size of 6 for Salinas, 3 for the Indian Pines dataset, 10 for the Pavia University dataset, and 4 for the PRISMA dataset. The batch size was selected based on the memory constraints of the GPUs and the dataset characteristics, ensuring stable and efficient training. The learning rate was set to a constant value of 0.01 across all experiments, as introducing a dynamic learning rate provided no significant benefits; the loss function consistently reached a plateau within the chosen number of epochs.
The model was trained for 30 epochs on the Salinas and Pavia University datasets, 50 epochs on the Indian Pines and PRISMA datasets, as these values ensured the loss function stabilized while avoiding overfitting. In all experiments, we used the Categorical Cross Entropy loss function to optimize performance.
To evaluate segmentation performance, we report metrics including Overall Accuracy (OA), Kappa Coefficient (Kappa), and Average Accuracy (AA). These metrics will be explained in detail in the following section.
@@ 4.4 Data Preprocessing
We constructed the datasets using only random patches, in contrast to other studies that employ various data balancing strategies (e.g., [26]). This approach aimed to evaluate AMBER’s ability to learn from unbalanced datasets and to create datasets with a consistent number of patches. The center of each patch was randomly selected from pixels that were not located on the edges of the images. For the Salinas, Indian Pines, and Pavia University datasets, the center was chosen from pixels that did not belong to the undefined class, ensuring patches contained meaningful information.
To mitigate the well-known issue of pixel overlap between training and test sets, which can lead to information leakage [27, 28], we allocated a small percentage of each dataset for training. This methodology aligns with established practices in the literature (e.g, [5, 29, 30]). Specifically, for the Salinas, Indian Pines and Pavia University datasets, 20% of the patches were assigned to the training set, while the remaining 80% were used for testing. For the PRISMA dataset, the training set comprised only 10% of the patches, with 90% reserved for testing.
Before the training phase, dimensionality reduction techniques, such as PCA, SVD, or Gabor filters [29, 13, 14], are commonly applied to the spectral dimension of hyperspectral images to address information redundancy. However, in our approach, we retained the full spectral dimension of the patches. This decision aimed to enable fine-grained classification and delegate the task of extracting relevant features to AMBER’s encoder.
@@ 5.1 Classification Results
To assess the effectiveness of the proposed method in classifying hyperspectral remote sensing images, we compared it against four mainstream approaches: SVD/Unet [14], HSI-CNN [31], 3D-CNN [32], and contextual deep CNN [33]. Notably, not all methods employ pixel-by-pixel classification. Among these methods, only SVD/Unet, contextual deep CNN, and AMBER utilize this approach, whereas HSI-CNN and 3D-CNN focus on image-level classification, where class labels are assigned based on the central pixel of each patch. This distinction inherently boosts the reported metrics for the latter two methods. Despite this, AMBER achieves superior performance on the Salinas, Indian Pines and Pavia University datasets and delivers results comparable to the best methods on the PRISMA dataset.
@@
The3D-CNNarchitecture is a Multiscale 3D deep convolutional neural network (M3DDCNN) which could jointly learn both 2D Multi-scale spatial feature and 1D spectral feature from HSI data in an end-to-end approach.
@@
Each neural network mentioned above was trained on the same patches using the network settings and configurations described in the corresponding papers [14, 31, 32, 33]. In this context, “same patches” refers to patches centered on the same pixel. AMBER was trained on patches of dimension $D\times 32\times 32$, where $D$ represents the spectral dimension, as previously mentioned. SVD/Unet was trained on patches of $3\times 32\times 32$ (3 corresponding to the top three singular values), HSI-CNN on patches of $D\times 3\times 3$, 3D-CNN on patches of $D\times 7\times 7$, and contextual deep CNN on patches of $D\times 5\times 5$. As described in most articles (e.g., [29, 30, 34, 5]), the datasets were divided into training and testing sets by random sampling throughout the image, ensuring that the samples did not overlap. In this context, “samples” refer to the central pixel of the patch.
For all methods, we conducted five independent experiments on each dataset. In each experiment, we used five-fold Monte Carlo cross-validation, using various random sampling of the entire image to generate different training and test sets [35]. This approach ensured robustness and reliability in evaluating the model’s performance.
The results presented in Tables 5, 6, 7, and 8 represent the mean classification accuracy for each class, along with the mean Overall Accuracy (OA), Kappa coefficient, and Average Accuracy (AA) computed across the five-fold Monte Carlo cross-validation sets, together with their standard deviations (±) to illustrate the variability and reliability of the model’s performance across different validation folds. The highest values for each metric are highlighted in bold.
Table 5: Classification accuracy of Salinas Dataset.
Class SVD/Unet [14] HSI-CNN [31] 3D-CNN [32] Cont Deep CNN [33] AMBER
1 $0.00$ $99.10$ $99.49$ $97.14$ $\boldsymbol{99.95}$
2 $99.99$ $97.29$ $99.99$ $99.82$ $\boldsymbol{100}$
3 $34.24$ $84.52$ $\boldsymbol{99.97}$ $99.31$ $99.65$
4 $91.47$ $96.91$ $98.60$ $99.52$ $\boldsymbol{99.68}$
5 $99.55$ $98.49$ $98.09$ $94.17$ $\boldsymbol{99.65}$
6 $99.96$ $99.98$ $\boldsymbol{100}$ $99.99$ $99.99$
7 $65.26$ $99.69$ $99.96$ $\boldsymbol{99.99}$ $99.98$
8 $97.96$ $86.49$ $96.09$ $90.95$ $\boldsymbol{99.46}$
9 $84.33$ $99.12$ $99.85$ $\boldsymbol{99.95}$ $96.99$
10 $79.15$ $91.84$ $96.76$ $95.63$ $\boldsymbol{99.74}$
11 $31.48$ $96.92$ $98.91$ $95.98$ $\boldsymbol{99.63}$
12 $32.15$ $\boldsymbol{100}$ $99.37$ $99.92$ 99.9399.93{99.93}99.93
13 $99.57$ $98.90$ $95.29$ $\boldsymbol{99.67}$ $99.57$
14 $99.05$ $95.56$ $\boldsymbol{99.83}$ $95.52$ $99.16$
15 $0.013$ $56.94$ $77.89$ $78.32$ $\boldsymbol{97.60}$
16 $64.54$ $96.34$ $99.11$ $99.16$ $\boldsymbol{99.97}$
OA/$72.28\pm 2.94$ $90.20\pm 0.65$ $96.08\pm 0.36$ $84.42\pm 2.02$ $\boldsymbol{99.73\pm 0.053}$
Kappa × 100 $67.87\pm 3.42$ $88.99\pm 0.029$ $95.59\pm 0.42$ $81.05\pm 2.56$ $\boldsymbol{99.69\pm 0.061}$
AA/$67.42\pm 6.80$ $93.63\pm 0.24$ $97.46\pm 0.19$ $71.40\pm 9.02$ $\boldsymbol{99.75\pm 0.021}$
[…]
@@
Figure 6: Salinas prediction. (a) False-color map. (b) Ground-truth map. (c) AMBER prediction with the undefined mask. (d) AMBER prediction without the undefined mask.
@@
The AMBER method consistently outperformed the compared methods on the Salinas, Indian Pines, and Pavia University datasets, achieving the highest Overall Accuracy (OA), Kappa coefficients (Kappa) and Average Accuracy (AA) across five independent experiments.
For the Salinas dataset (Table 5), AMBER achieved the best OA of $99.73\%\pm 0.053$, the best Kappa coefficient of $99.69\pm 0.061$, and the best AA of $99.75\%\pm 0.021$. These metrics surpass all competing methods, including 3D-CNN ($96.08\%\pm 0.36$ OA) and HSI-CNN ($90.20\%\pm 0.65$ OA). AMBER outperformed other networks in 10 out of the 16 classes, achieving near-perfect accuracy in classes such as Class 2 ($100\%$) and Class 6 ($99.99\%$). Furthermore, AMBER demonstrated its robustness in minority classes like Class 15, achieving $97.60\%$, significantly higher than HSI-CNN ($56.94\%$) and SVD/Unet ($0.013\%$).
For the Indian Pines dataset (Table 6), AMBER again outperformed all other models, achieving an OA of $99.74\%\pm 0.097$, Kappa coefficient of $99.68\pm 0.12$, and AA of $99.18\%\pm 0.34$. AMBER demonstrated exceptional accuracy across underrepresented classes, such as Class 7 (Grass-pasture-mowed), where it achieved $96.02\%$, while other methods like HSI-CNN and SVD/Unet failed entirely, with $0\%$ accuracy. This ability to handle class imbalance highlights AMBER’s robustness. AMBER also achieved the highest or near-highest accuracy in most other classes, such as Class 13 (Soybean-mintill) and Class 14 (Soybean-clean), where it delivered $99.95\%$ and $99.97\%$, respectively.
For the Pavia University dataset (Table 7), AMBER achieved the best OA ($99.94\%\pm 0.0098$), Kappa coefficient ($99.91\pm 0.015$), and AA ($99.82\%\pm 0.022$). It consistently delivered the highest accuracy across most classes, including minority classes like Class 3 (Gravel), where it achieved $99.27\%$, significantly outperforming SVD/Unet ($0.17\%$). AMBER also achieved near-perfect accuracy in dominant classes, such as Class 2 (Asphalt) and Class 9 (Shadows), achieving $100\%$ and $99.98\%$, respectively. Additionally, its low standard deviations across all metrics underscore its reliability and generalization capabilities.
For the PRISMA dataset (Table 8), AMBER achieved an OA of $90.90\%$, a Kappa coefficient of $0.8741$, and an AA of $90.92\%$. While these results were slightly lower than the SVD/Unet method, AMBER still outperformed other competing approaches. The marginally lower performance is under investigation, suggesting potential areas for further model optimization.
AMBER demonstrates exceptional performance across hyperspectral datasets, outperforming state-of-the-art methods such as 3D-CNN, HSI-CNN, and SVD/Unet on the Salinas, Indian Pines, and Pavia University datasets. It consistently achieves high accuracy in both majority and minority classes, effectively classifying challenging features such as Broccoli-green weeds (Salinas), Alfalfa, Grass-pasture-mowed, and Soybean-mintill (Indian Pines), and Asphalt, Gravel, and Bitumen (Pavia University). On the PRISMA dataset, AMBER achieves state-of-the-art performance, successfully identifying complex features such as mining areas and built-up regions. Its ability to generalize across both airborne datasets (Salinas, Indian Pines, Pavia University) and spaceborne data (PRISMA) is evident from its consistently high accuracy and low standard deviations in OA, Kappa and AA, underscoring its robustness and reliability.
@@
The Fig. 6, 7, 8, and 9 provide visual results for each dataset, including the false-color image (a), ground-truth sample plots (b), and AMBER’s classification outputs (c) and (d). In particular, plot (c) shows the predicted classification results, while plot (d) overlays a mask of ground-truth undefined pixels to visually justify AMBER’s high accuracy across all feature types presented in Tables 5, 6, 7, and 8. Notably, for the Salinas, Pavia University, and Indian Pines datasets, the ground-truth map (b) and AMBER’s predictions (c) are nearly identical.
@@
Figure 10 highlights AMBER’s ability to accurately predict pixels classified as undefined in the ground-truth map, which are excluded from the loss function computation. The figure illustrates AMBER’s capacity to segment complex structures. The example shows an asphalt class structure in the Pavia University dataset.